# Natural solution of SUSY $\mu$ problem from modulus stabilization in modular flavor model

### v3

@@ 6 Conclusions
In this study, we propose a novel solution to the SUSY $\mu$-problem within the framework of modular flavor symmetry. The explicit $\mu$-term is prohibited by modular symmetry after assigning non-vanishing modular weights and/or non-trivial singlet representations to the $H_{u},H_{d}$ superfields. The effective $\mu$-term is then regenerated following the stabilization of the modulus field, with the corresponding modulus VEV breaks the modular symmetry partially (at the fixed points) or completely (in the vicinity of the fixed points). Since naturalness requires $\mu_{eff}$ to lie at the electroweak scale or soft SUSY breaking scale, a large suppression factor leading to $\mu_{eff}\ll\Lambda_{SUSY}$ is required when the modulus field is stabilized. Such a large suppression factor can naturally arise from a modulus VEV in the vicinity of certain fixed points.
Ordinary single modulus stabilization mechanism can be affected by SUSY breaking contributions. The Goldstino, which is necessarily present as a consequence of UV SUSY breaking and mixes with gravitino, can be promoted into the constrained nilpotent superfield $\textbf{X}_{nl}$ and provides a way for an arbitrary non-supersymmetric low-energy theory to realize SUSY non-linearly. We discuss the stabilization mechanism of a single modulus field in the presence of SUSY breaking contributions, described by the non-linear SUSY realization scheme with a nilpotent Goldstino $\textbf{X}_{nl}$. The presence of $\textbf{X}_{nl}$ contributions can slightly alter the original stabilized modulus VEV to a value near the fixed points.
@@
The modulus stabilization mechanism with nilpotent $\textbf{X}_{nl}$ superfield can be fairly generic and ubiquitous. The regeneration of small $\mu_{eff}$ term following modulus stabilization, which just needs a modulus VEV near some fixed points, is independent on the concrete modulus stabilization mechanism.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China under grant numbers Nos.12075213 and 12335005, by the Natural Science Foundation for Distinguished Young Scholars of Henan Province under grant number 242300421046.

### v4

@@ 6 Conclusions
In this study, we propose a novel solution to the SUSY $\mu$-problem within the framework of modular flavor symmetry. The explicit $\mu$-term is prohibited by modular symmetry after assigning non-vanishing modular weights and/or non-trivial singlet representations to the $H_{u},H_{d}$ superfields. The effective $\mu$-term is then regenerated following the stabilization of the modulus field, with the corresponding modulus VEV breaks the modular symmetry partially (at the fixed points) or completely (in the vicinity of the fixed points). Since naturalness requires $\mu_{eff}$ to lie at the electroweak scale or soft SUSY breaking scale, a large suppression factor leading to $\mu_{eff}\ll\Lambda_{SUSY}$ is required when the modulus field is stabilized. Such a large suppression factor can naturally arise from a modulus VEV in the vicinity of certain fixed points or by the suppression factor in weighton-like mechanism.
Ordinary single modulus stabilization mechanism can be affected by SUSY breaking contributions. The Goldstino, which is necessarily present as a consequence of UV SUSY breaking and mixes with gravitino, can be promoted into the constrained nilpotent superfield $\textbf{X}_{nl}$ and provides a way for an arbitrary non-supersymmetric low-energy theory to realize SUSY non-linearly. We discuss the stabilization mechanism of a single modulus field in the presence of SUSY breaking contributions, described by the non-linear SUSY realization scheme with a nilpotent Goldstino $\textbf{X}_{nl}$.
@@
The modulus stabilization mechanism with nilpotent $\textbf{X}_{nl}$ superfield can be fairly generic and ubiquitous. The regeneration of small $\mu_{eff}$ term following modulus stabilization, which just needs a modulus VEV near some fixed points or through a weighton-like mechanism for $H_{u}H_{d}$ bilinear, is independent of the concrete modulus stabilization mechanism.
@@ Appendix A Properties of Modular $A_{4}$
Appendix B Eisenstein Series
The Eisenstein series $G_{k}(\tau)$, defined by
[…]
is a modular form of weight $k$ for $\tau\in{\cal H}$ and has the following Fourier decomposition
$G_{k}(\tau)=2\zeta(k)+2\frac{(2\pi i)^{k}}{\Gamma(k)}\sum\limits_{N=1}^{\infty}\sigma_{k-1}(N)q^{N}~,$ (133)
with $\sigma_{\alpha}(N)$ the sums of divisors functions, defined for any $\alpha\in\mathcal{C}$ by
$\sigma_{\alpha}(N)=\sum\limits_{l|N}l^{\alpha}~.$ (134)
It can be proven that $G_{4}(\rho)=G_{6}(i)=0$.
Factoring out $2\zeta(k)$, and using the relation between Bernoulli numbers $B_{k}$ and $\zeta(k)$ for even $k$, the normalized Eisenstein series $E_{k}$ can be defined by
$E_{k}(\tau)=\frac{G_{k}(\tau)}{2\zeta(k)}=1-\frac{2k}{B_{k}}\sum\limits_{N=1}^{\infty}\sigma_{k-1}(N)q^{N}~,$ (135)
with some of the $\nu_{k}\equiv-{2k}/{B_{k}}$ values given by
$\nu_{4}=240,~\nu_{6}=504,~\nu_{8}=480.$ (136)
The quasi-modular transformation property of $E_{2}(\tau)$
$E_{2}(\gamma\tau)$ $=$ $(c\tau+d)^{2}E_{2}(\tau)+\frac{12}{2\pi i}c(c\tau+d)~,$ (137)
allows one to evaluate $E_{2}(\tau)$ at the points $\tau=i$ and $\tau=\omega$.
The values of $E_{k}(\tau)$ are listed below DHoker:2022dxx
$E_{2}(i)$ $=$ $\frac{3}{\pi},~~E_{2}(\omega)=\frac{2\sqrt{3}}{\pi},~~E_{4}(i)=\frac{48[\Gamma(\frac{5}{4})]^{4}}{\pi^{2}[\Gamma(\frac{3}{4})]^{4}},$~
$E_{4}(\omega)$ $=$ $0,~~E_{6}(i)=0,~~E_{6}(\omega)=\frac{729}{2\pi^{3}}\frac{[\Gamma(\frac{4}{3})]^{6}}{[\Gamma(\frac{5}{6})]^{6}}.$ (138)
The second derivative of $j(\tau)$ are given by
$j^{\prime\prime}(\tau)=\frac{4\pi^{2}}{[\eta(\tau)]^{24}}\left\{\frac{1}{6}E_{2}(\tau)[E_{4}(\tau)]^{2}E_{6}(\tau)-\frac{2}{3}E_{4}(\tau)[E_{6}(\tau)]^{2}-\frac{1}{2}[E_{4}(\tau)]^{4}\right\}~,$ (139)
from Ramanujan’s formulas for derivatives of Eisenstein series, which gives
$j^{\prime\prime}(\omega)=0~,~~j^{\prime\prime}(i)=-\frac{2\pi^{2}}{[\eta(i)]^{24}}[E_{4}(i)]^{4}~.$ (140)
Appendix C Solutions of $H^{\prime}(\tau)=0$
Such a constraint reduces to
$H^{\prime}(\tau)=\left\{j(\tau)^{n/3}[j(\tau)-1728]^{m/2}\mathcal{P}(j(\tau))\right\}^{\prime}=0,$ (141)
or equivalently,
$H(\tau)j^{\prime}(\tau)\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{\mathcal{P}(j(\tau))}\frac{\partial\mathcal{P}(j(\tau))}{\partial j(\tau)}\right]=0.$ (142)
Note that the expression
$H(\tau)\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{\mathcal{P}(j(\tau))}\frac{\partial\mathcal{P}(j(\tau))}{\partial j(\tau)}\right],$ (143)
can be singular at certain solutions of $H(\tau)=0$. Such singular behavior would be eliminated by the multiplication of $j^{\prime}(\tau)$ to render $H^{\prime}(\tau)$ finite at such points, which can be seen in (45).
Firstly, we would like to discuss the case $H(\tau)\neq 0$. From equation (143), it is obvious that either
$\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{\mathcal{P}(j(\tau))}\frac{\partial\mathcal{P}(j(\tau))}{\partial j(\tau)}\right]=0~,$ (144)
or $j^{\prime}(\tau)=0$.
When $H(\tau)\neq 0$, for non-negative integers $m$ and $n$ (by the definition of $H(\tau)$), we have the following discussions for equation (144) after eliminating the $H(\tau)$ factor:
  • $m+n>0$:
    • When $n\neq 0$ and $m=0$, condition (144) reduces to
      $\frac{d\mathcal{P}(j(\tau))}{\mathcal{P}(j(\tau))}=-\frac{n}{3}\frac{dj(\tau)}{j(\tau)},$ (145)
      which gives
      $\mathcal{P}(j(\tau))=C_{0}[j(\tau)]^{-n/3}.$ (146)
      So, $\mathcal{P}(j(\tau))$ is not a polynomial of $j(\tau)$ unless $n\in 3\mathbb{Z}$ is negative, contradicting the assumption that $n$ is non-negative integers.
    • When $m\neq 0$ and $n=0$, it can also be deduced that $\mathcal{P}(j(\tau))$ is a polynomial of $j(\tau)$ unless $m\in 2\mathbb{Z}$ is negative, contradicting the assumption that $m$ is non-negative integers.
    • For $m\neq 0$ and $n\neq 0$, similar deduction gives
      $\mathcal{P}(j(\tau))=C_{0}^{\prime}[j(\tau)]^{-n/3}[j(\tau)-1728]^{-m/2}.$ (147)
      So, $\mathcal{P}(j(\tau))$ is not a polynomial of $j(\tau)$ unless both $m\in 2\mathbb{Z}$ and $n\in 3\mathbb{Z}$ are negative. As $m,n$ are defined to be non-negative integers, this condition requires $m=n=0$, contradicting the assumptions that $m,n\neq 0$.
    Therefore, $m+n>0$ can not hold.
  • $m=0$ and $n=0$:
    With the constraints $m=n=0$, the condition (143) reduces to
    $\frac{\partial\mathcal{P}(j(\tau))}{\partial j(\tau)}=0,$ (148)
    which gives
    $\mathcal{P}(j(\tau))=\text{Constant}.$ (149)
    Since $H(\tau)\neq 0$, this constant must therefore be non-zero.
When $H(\tau)\neq 0$, for the condition $j^{\prime}(\tau)=0$, the derivative of the Klein $j$-function is given by
$\frac{\partial}{\partial\tau}j(\tau)=-2\pi i\frac{{E_{6}(\tau)}{E_{4}^{2}(\tau)}}{[\eta(\tau)]^{24}}=0~.$ (150)
It can checked that its solutions are $\omega$ and $i$, contradicting the assumption $H(\tau)\neq 0$ with $m=n=0$. Therefore, the condition $H^{\prime}(\tau)=0$ with $H(\tau)\neq 0$ implies that $H(\tau)$ is a non-zero constant.
The solutions of $H^{\prime}(\tau)=0$ for the case $H(\tau)=0$ are addressed in the paragraphs containing equations (70), (71), and (72).
@@ Acknowledgements.
These authors contributed equally: Hong Jie Fan, Ying Kai Zhang. We are very grateful to the referees for helpful discussions. This work was supported by the National Natural Science Foundation of China under grant numbers Nos.12075213 and 12335005, by the Natural Science Foundation of Henan (Distinguished Young Scholars of Henan Province) under grant number 242300421046.